after promotion, Gewn's annual salary was increased by 7.85% to R 450000. Determine his previous monthly salary and the monthly increase he will be getting


1
Expert's answer
2022-03-07T11:10:32-0500

"R 450000= 107.85%"

? = 100

"\frac{450000\times 100}{107.85}" = 417246.1752433936

Monthly Salary= "\frac {417246.1752433936 }{12}"

= 34, 770.51

37,500- 34, 770.51= 2,729.49


Increment per month= 227.4575